# Convergence and Optimality of the EM Algorithm Under Multi-Component Gaussian Mixture Models

### v1

@@ D.2.3 Proof of Lemma 25 for $j=0$
for some constant $0<\nu<(1/2-2c_{0}-2c_{0}^{2})$. Further let $S(\omega)]^{c}:=[n]\setminus S(\omega)$ be its complement set. We first observe that for any $i\in\widehat{G}_{\ell}\cap[S(\omega)]^{c}$, using (C.56), Lemma 21 and the choice of $\nu$ gives
@@ D.3 Proof of Theorem 8 & Theorem 9
The last two displays ensure that $\widehat{\theta}^{(t)}$ satisfies (3.5) hence $\widehat{\omega}^{(t)}\in\mathcal{B}_{d}(\omega^{*})$. By Sections D.3 and D.53 and induction, we obtain
@@ D.4 Proof of Remark 4
Bing et al. [7, Lemmas 16 & 17] give that for any $t\geq 0$, with probability at least $1-4e^{-t}$,
@@
By inspecting the proof of Bing et al. [7, Lemma 15], one can also deduce that
@@ E.1 Proof of Theorem 11
According to Lemmas 33, 32 and 36, the intersection event holds with probability $1-e^{-n\pi_{\min}^{*}}-n^{-2}-n^{-1}\geq 1-2n^{-1}$ by $n\pi_{\min}^{*}\geq C\log n$. The event $\mathcal{E}_{\ell_{2}}$ ensures that
@@
by $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$ and (4.15). On the other hand, when $\widehat{\theta}^{(t)}$ meets the requirement in (b), then Eqs. E.27, E.28 and E.29 and (4.16) ensure that $\widehat{\theta}^{(t)}$ also satisfies (3.5). The proof is complete. ∎
@@ Appendix F Proof of Theorem 14: the minimax lower bounds
which jointly imply the claim. We first note that when $d<16$, the lower bounds of $c/\sqrt{n\alpha}$ and $c/\sqrt{n}$ can be easily proved by applying Le Cam’s two-point argument; see, Tsybakov [40, Theorem 2.2 & Eq (2.9)]. We thus focus on $d\geq 16$ in the following.
@@
Note that the existence of such $\mu_{\ell}$’s is guaranteed by $d\geq L-1$. For all $0\leq i\leq N$, let $\theta^{(i)}=(\pi,\mu_{1}+\varepsilon\beta^{(i)},\mu_{2},\ldots,\mu_{L},\mathbf{I}_{d})$ with $\beta^{(0)}=0_{d}$ and $\beta^{(i)}\in\{0,1\}^{d}$ for $i\geq 1$, which according to the Varshamov-Gilbert bound in Tsybakov [40, Lemma 2.9], satisfy
@@
We proceed to invoke Tsybakov [40, Theorem 2.5] to prove (F.1), which requires to verify
@@
invoking Tsybakov [40, Theorem 2.5] proves (F.1) for $c_{1}=\sqrt{c\log 2}/16$ and
@@
In view of (F), choosing $c$ sufficiently small, invoking Tsybakov [40, Theorem 2.5] proves (F.7) for $c_{1}=\tau/(12\sqrt{2})$ and $d\geq 16$. This completes the proof. ∎
@@ Appendix G Auxiliary Lemma
The following lemmas are proved in [8].
@@
The following lemma states upper bounds of the quadratic form of a sub-Gaussian random vector [24].
@@
The following lemmas state the well-known vector-valued and matrix-valued Bernstein inequalities. See, for instance, Minsker [33, Theorem 3.1, Corollary 3.1 and Corollary 4.1].
@@
Another useful concentration inequality of the operator norm of the random matrices with i.i.d. sub-Gaussian rows is stated in the following lemma [6, Lemma 16]. This is an immediate result of Vershynin [42, Remark 5.40].

### v2

@@ D.2.3 Proof of Lemma 24 for $j=0$
for some constant $0<\nu<(1/2-2c_{0}-2c_{0}^{2})$. Further let $S(\omega)]^{c}:=[n]\setminus S(\omega)$ be its complement set.
Recall the model (1.1), we partition $n]=\bigcup_{a=1}^{L}\widehat{G}_{a}$ with
Further let
We first observe that for any $i\in\widehat{G}_{\ell}\cap[S(\omega)]^{c}$, using (C.54), Lemma 20 and the choice of $\nu$ gives
@@ D.3 Proof of Theorem 7 & Theorem 8
The last two displays ensure that $\widehat{\theta}^{(t)}$ satisfies (3.5) hence $\widehat{\omega}^{(t)}\in\mathcal{B}_{d}(\omega^{*})$. By Sections D.3 and D.40 and induction, we obtain
@@ D.4 Proof of Remark 4
bing2024lineardiscriminantregularizedregression give that for any $t\geq 0$, with probability at least $1-4e^{-t}$,
@@
By inspecting the proof of bing2024lineardiscriminantregularizedregression, one can also deduce that
@@ E.1 Proof of Theorem 10
According to Lemmas 32, 31 and 35, the intersection event holds with probability $1-e^{-n\pi_{\min}^{*}}-n^{-2}-n^{-1}\geq 1-2n^{-1}$ by $n\pi_{\min}^{*}\geq C\log n$. The event $\mathcal{E}_{\ell_{2}}$ ensures that
@@
by $\Delta_{\min}\geq C\log(1/\pi_{\min}^{*})$ and (a). On the other hand, when $\widehat{\theta}^{(t)}$ meets the requirement in (b), then Eqs. E.27, E.28 and E.29 and (b) ensure that $\widehat{\theta}^{(t)}$ also satisfies (3.5). The proof is complete. ∎
@@ Appendix F Proof of Theorem 13: the minimax lower bounds
which jointly imply the claim. We first note that when $d<16$, the lower bounds of $c/\sqrt{n\alpha}$ and $c/\sqrt{n}$ can be easily proved by applying Le Cam’s two-point argument; see, Tsybakov09. We thus focus on $d\geq 16$ in the following.
@@
Note that the existence of such $\mu_{\ell}$’s is guaranteed by $d\geq L-1$. For all $0\leq i\leq N$, let $\theta^{(i)}=(\pi,\mu_{1}+\varepsilon\beta^{(i)},\mu_{2},\ldots,\mu_{L},\mathbf{I}_{d})$ with $\beta^{(0)}=0_{d}$ and $\beta^{(i)}\in\{0,1\}^{d}$ for $i\geq 1$, which according to the Varshamov-Gilbert bound in Tsybakov09, satisfy
@@
We proceed to invoke Tsybakov09 to prove (F.1), which requires to verify
@@
invoking Tsybakov09 proves (F.1) for $c_{1}=\sqrt{c\log 2}/16$ and
@@
In view of (F), choosing $c$ sufficiently small, invoking Tsybakov09 proves (F.7) for $c_{1}=\tau/(12\sqrt{2})$ and $d\geq 16$. This completes the proof. ∎
@@ Appendix G Auxiliary Lemma
The following lemmas are proved in [bing2023optimal].
@@
The following lemma states upper bounds of the quadratic form of a sub-Gaussian random vector [hsu2012tail].
@@
The following lemmas state the well-known vector-valued and matrix-valued Bernstein inequalities. See, for instance, Minsker2017.
@@
Another useful concentration inequality of the operator norm of the random matrices with i.i.d. sub-Gaussian rows is stated in the following lemma [bing2020prediction, Lemma 16]. This is an immediate result of vershynin_2012.